# Reconstructing Intelligible Speech from the Pressure Sensor Data in HVACs

### v1

@@ 6.6 Complex Multiresolution STFT Loss
In this work, we use Complex Multi-Resolution STFT Loss that measures the differences between clean and enhanced signals in the complex T-F domain (i.e., amplitude and phase). Unlike standard magnitude-only loss [55], we propose a complex multiresolution STFT loss that separately evaluates the loss on real and imaginary parts of STFT across multiple resolutions, thereby capturing fine and coarse spectral details in complex domains. At first, spectral convergence loss $L_{SC}$ [56] and log STFT magnitude loss $L_{mag}$ [56] are calculated on both real and imaginary parts of the enhanced signal and ground truth speech data. Let us define the $L_{SC}$ and $L_{mag}$ calculated on real and imaginary STFT data as {$L^{r}_{SC}$, $L^{i}_{SC}$} and {$L^{r}_{mag}$, $L^{i}_{mag}$}, respectively. Assuming we have $S$ different STFT resolutions, we aggregate the losses by averaging over the resolutions for both real and imaginary parts following Eqn. 9.
$L_{\mathrm{r}}=\frac{1}{S}\sum_{s=1}^{S}\Big{(}L_{\mathrm{SC}}^{r}+L_{\mathrm{mag}}^{r}\Big{)}$ (9)
$L_{\mathrm{i}}=\frac{1}{S}\sum_{s=1}^{S}\Big{(}L_{\mathrm{SC}}^{i}+L_{\mathrm{mag}}^{i}\Big{)}$
@@
We use three resolutions, such as frequency bins = [256, 512, 1024], hop sizes = [128, 256, 512], and window lengths = [256, 512, 1024] to calculate $L_{r}$ and $L_{i}$. The overall complex multiresolution STFT loss, $L^{complex}_{STFT}$ is the sum of $L_{r}$ and $L_{i}$. The joint optimization in complex T-F domain on magnitue and phase improves the quality and intelligibility of the reconstructed speech from pressure sensor data.
500 Hz to 8 kHz 1 kHz to 8 kHz 2 kHz to 8 kHz
@@
Raw pressure data 3.45 0.84 6.24 0.87 0.71 3.15 0.95 8.54 0.98 0.75 2.95 1.27 9.87 1.14 0.79
Reconstructed 1.24 1.78 8.78 1.51 0.75 1.16 1.95 10.04 1.61 0.79 1.01 2.17 11.38 1.95 0.82
TABLE IV: Evaluating WaLi for reconstructing intelligible audio from pressure sensor data for 500 Hz, 1 kHz, and 2 kHz sampling frequencies to 8 KHz upsampling for 70 dB audio. Here, L = LSD, N = NISQA-MOS, S = SI-SDR, P = PESQ, and ST = STOI.
7 Evaluation
7.1 Pressure Data and Audio Corpus
@@
We use VCTK (version 0.92) [57], a multi-speaker English corpus for evaluation. The dataset contains $\sim$43600 audio traces sampled at 48 kHz from 110 individuals (i.e., balanced for males and females). We downsample the dataset to 8 kHz for evaluation. Each audio clip has a duration ranging from 2s to 7s. We standardize all audio clips to 4s by either zero-padding or silence trimming. Following [58], only the mic1 data are used for the experiments, and p280 and p315 are omitted due to technical issues. We play the audio at 60 dB through a loudspeaker and collect the corresponding pressure data via a DPS sensor (SDP810-125PA). The speaker is placed at a 5 cm distance from one of the pressure ports to play the audio clips. To prevent sound signals from influencing the pressure in the other port, we connect a sampling tube to it, isolated by positioning its opening a meter away. Signals from the sensor are collected as .edf and next converted to .wav format using a custom Python script. The sampling frequency of the pressure sensor is varied in between 500 Hz and 2 kHz. We use sinc interpolation to upsample before the audio reconstruction to ensure that the system input and output have the same shape. Note that in a real case, the speech contents may be different from the spoken ones during the attack phase. Thus, for testing purposes, we use 11 different speakers, who are not present in the training. The models are trained offline with an NVIDIA 4090 GPU and deployed on a desktop for speech recovery.
7.2 Evaluation Metrics
We comprehensively evaluate WaLi using the following five different metrics.
b) Non-Intrusive Speech Quality Assessment - Mean Opinion Score (NISQA-MOS): NISQA-MOS is used to estimate the perceived quality of reconstructed audio without needing a reference signal. It predicts a MOS similar to human ratings with values between 0 to 5.
e) Short-Time Objective Intelligibility (STOI): STOI is specifically designed to predict how understandable speech is to human listeners, especially in the presence of noise. STOI has values between 0 to 1.
7.3 Overall Performance
@@
To intuitively observe the performance of WaLi, we compared among ground truth speech, raw data from the pressure sensor, and reconstructed speech by WaLi, shown in Fig. 8. We can see that the frequency components above 250 Hz are absent in the pressure sensor data (see Fig. 8(b)) as we use a sampling frequency of 0.5 kHz in DPS at this time. The ground truth audio is sampled at 8 kHz (4 kHz bandwidth), shown in Fig. 8(a). From Fig. 8(c), we can see that the high-frequency components are reconstructed up to 4 kHz by WaLi. The reconstructed spectrogram shows a high similarity to the ground truth one.
@@
Moreover, to quantitatively measure WaLi’s performance, we vary the sampling frequency of the pressure sensor on three scales, such as 500 Hz, 1 kHz, and 2 kHz. We reconstruct the audio to 8 kHz upsampling, to evaluate WaLi for high upsampling ratios [55]. The detailed comparison is shown in Table IV, which shows the average value of the metrics over all the test speakers. The reconstructed audio by WaLi achieves a 2.78x improvement in LSD (i.e., 3.45 vs 1.24), 2.11x increase in NISQA-MOS (i.e., 0.84 vs 1.78), 1.4x increase in SI-SDR (i.e., 6.24 vs 8.78), 1.73x increase in PESQ (i.e., 0.87 vs 1.51), and 1.05x increase in STOI (i.e., 0.71 vs 0.75) between the raw pressure data and the reconstructed audio for 500 Hz sampling frequency.
@@
Please also note that the metrics improve for 2-8 kHz compared to 1-8 kHz and 0.5-8 kHz because for 2-8 kHz the upsampling ratio is around 4, for 1-8 kHz the upsampling ratio is 8, and for 0.5-8 kHz the upsampling ratio is 16. The lower the upsampling ratio, the greater the performance gain for the reconstruction model.
Individual test speaker: In addition to average values, Fig. 9 (Left) shows the individual values of each test speaker for LSD and NISQA-MOS. The individual values indicate that WaLi performs consistently for all unseen test speakers. This indicates that WaLi can effectively reconstruct the high-frequency components and suppress artifacts, and improve the speech intelligibility significantly.
@@
Figure 9: (Left) LSD and NISQA-MOS for individual test speakers. (Right) Comparison between WaLi and VibSpeech [10] in terms of MCD and SNR.
Comparison with VibSpeech [10]: We also compare between WaLi and VibSpeech [10] in terms of SNR and Mel-Ceptral Distortion (MCD) [10] in Fig. 9 (Right), which indicates that WaLi performs better (MCD = 3.81 and SNR = 6.1) than VibSpeech (MCD = 3.9 and SNR = 5.4).
Please note that WaLi is tested with different speakers, which are not seen by the model during the training time. Moreover, the VCTK dataset consists of everyday-use audio clips, which are not limited to a certain specialized vocabulary such as [34]. Therefore, WaLi can recover intelligible audio with unrestricted vocabulary irrespective of specific applications.
7.4 Phase Reconstruction in Non-Noisy Condition
As WaLi is a complex-valued model, it can directly reconstruct phase information from the low-resolution pressure sensor data. Previous works used the Griffin-Lim algorithm [13] as a vocoder to reconstruct the phase. To compare the performance of our phase reconstruction in the non-noisy condition, we reconstruct audio from only the magnitude of our model with the Griffin-Lim algorithm, and then compare it with the reconstructed audio from our model. A comparison summary is shown in Table V for a 500 Hz to 8 kHz reconstruction.
@@ 7.4 Phase Reconstruction in Non-Noisy Condition
TABLE V: Evaluating WaLi for reconstructing phase at non-noisy condition. Here, L = LSD, N = NISQA-MOS, S = SI-SDR, P = PESQ, and ST = STOI.
Table V indicates that our proposed complex-valued WaLi is similar/slightly better in phase reconstruction compared to the Griffin-Lim algorithm in the non-noisy condition. However, the performance of WaLi is much better under noisy conditions that is discussed in the next section.
7.5 Phase Reconstruction in Noisy Conditions
DPSs, operating within a low-pressure range of 0-200 Pa and at high sampling frequencies of 0.5-2 kHz, are sensitive to transient noise sources (see Section 3.6). Therefore, clean phase reconstruction becomes particularly significant in noisy conditions. To test the performance of phase reconstruction under noisy conditions, we use the DCASE challenge dataset [59] to add transient mechanical noise, such as fans, valves, pump noise, air leaks, tool drops, etc., to the clean VCTK dataset. Noise is randomly added within a range of -7 to 40 dB. Then, we retrain WaLi with a {noisy, clean} pair dataset. A comparison summary is shown in Table VI for a 500 Hz to 8 kHz reconstruction in noisy conditions. Table VI indicates that our proposed complex-valued WaLi is much better in phase reconstruction compared to the Griffin-Lim algorithm in transient noisy condition. The reason behind this is that WaLi handles both magnitude and phase jointly and recovers both clean magnitude and phase under noisy conditions from the clean reference signal using a complex multi-resolution loss function. On the other hand, Griffin-Lim-based models [13] do not learn clean phase reconstruction as these models are magnitude-only models and cannot handle phase jointly.
@@ 7.5 Phase Reconstruction in Noisy Conditions
Griffin-Lim 1.38 1.43 6.17 1.21 0.70
WaLi 1.26 1.71 8.14 1.47 0.74
TABLE VI: Evaluating WaLi for reconstructing phase at transient noisy condition. Here, L = LSD, N = NISQA-MOS, S = SI-SDR, P = PESQ, and ST = STOI.
7.6 Impact of Sound Volume
We use a loudspeaker to simulate a speaker and vary the volume of the loudspeaker from 60 dB to 85 dB with a 5 dB increment. We inject the audio into the pressure sensor ports through the 1m long sampling tube from a 5 cm distance. The result is shown in Fig. 10 (Left) for LSD and NISQA-MOS for 500 Hz to 8 kHz upsampling. We can see that the LSD and NISQA-MOS improve rapidly between 60 dB and 70 dB. However, after 70 dB, LSD and NISQA-MOS improve steadily rather than a sharp improvement. The reason behind this is that the pressure sensor’s sensitivity increases with the increase of volume up to 70 dB. After 70 dB, the increase in volume does not drastically change the low-resolution frequency components. Therefore, the improvement is not sharp after 70 dB; rather, a steady ascent occurs. The LSD is close to 1, and NISQA-MOS is close to 2 for a wide volume range from normal conversation (i.e., 60 dB) to loud speech. It indicates that WaLi is capable to reconstruct intelligible audio from normal conversation (i.e., 60 dB) to loud speech in real-world scenario.

### v2

@@ IV-F Complex Multiresolution STFT Loss
Unlike standard magnitude-only loss [26], we propose a Complex Multi-Resolution STFT Loss that separately evaluates the loss on real (i.e., amplitude) and imaginary parts (i.e., phase) of STFT across multiple resolutions, thereby capturing fine and coarse spectral details in complex domains. At first, spectral convergence loss $L_{SC}$ [31] and log STFT magnitude loss $L_{mag}$ [31] are calculated on both real and imaginary parts of the enhanced signal and ground truth speech data. Let us define the $L_{SC}$ and $L_{mag}$ calculated on real and imaginary STFT data as $L^{r}_{SC}$, $L^{i}_{SC}$} and $L^{r}_{mag}$, $L^{i}_{mag}$}, respectively. Assuming we have $S$ different STFT resolutions, we aggregate the losses by averaging over the resolutions for both real and imaginary parts, following Eqn. 9.
$L_{\mathrm{r}}=\frac{1}{S}\sum_{s=1}^{S}\Big(L_{\mathrm{SC}}^{r}+L_{\mathrm{mag}}^{r}\Big);\,\,L_{\mathrm{i}}=\frac{1}{S}\sum_{s=1}^{S}\Big(L_{\mathrm{SC}}^{i}+L_{\mathrm{mag}}^{i}\Big)$ (9)
@@
We use $S$ = 3 resolutions, such as frequency bins = [256, 512, 1024], hop sizes = [128, 256, 512], and window lengths = [256, 512, 1024] to calculate $L_{r}$ and $L_{i}$. The overall complex multiresolution STFT loss, $L^{complex}_{STFT}$ is the sum of $L_{r}$ and $L_{i}$. The joint optimization in complex T-F domain on magnitude and phase improves the quality and intelligibility of the speech reconstructed from pressure sensor data.
V Attack Model Demonstration in Cleanroom
We demonstrate our attack at an FDA-compliant cleanroom located in an anonymous facility shown in Fig.7 (Left). The facility uses an industry-used DPS from Sensiron with part# SDP810-125PA [32]. It has two input ports connected to two vinyl sampling tubes with inner diameters of 3/16” and 5/16” [33]. A pressure pickup device with part# A-417A [34] is connected to one input port (see Fig. 1). A volunteer speaks from 0.5 m distance from the pressure pickup device. We record the output data from the DPS with a sampling frequency of 1 kHz. The recorded data from the DPS are shown in Fig. 7 (Middle) with its matching ground-truth audio (see WaLi).
In Fig. 7, the time domain waveform indicates a strong correlation between the ground truth audio and the pressure sensor data. We measure the intelligibility of the pressure sensor data by taking the ground truth as a reference using a metric named Perceptual Evaluation of Speech Quality (PESQ). PESQ is widely used to evaluate how a degraded speech signal compares to a reference (ground truth) signal, closely modeling human perception. The PESQ for the pressure data is 0.98. PESQ can have values between -0.5 to 4.5, where -0.5 means low perceptual quality and 4.5 means perceptually close to the ground truth. The 0.98 of PESQ for the pressure data indicates that the DPS captures low-grade perceptual audio, which is not intelligible (to be intelligible PESQ value of more than 1.4 is required). Therefore, it indicates that DPS can be used to eavesdrop by an attacker if the attacker can reconstruct intelligible audio from the severely aliased and low-grade perceptual data collected from DPSs.
Figure 7: (Left, Middle) Demonstration of attack on real-world setup. (Right) Experimental setup for attack model evaluation using BMS and DPSs.
TABLE II: Evaluating WaLi for reconstructing speech from pressure sensor data for 500 Hz, 1 kHz, and 2 kHz to 8 kHz upsampling for 60 dB audio. Here, L = LSD, N = NISQA-MOS, S = SI-SDR, P = PESQ, and ST = STOI, W = Whisper, and A = AssemblyAI.
Sensor 2 kHz to 8 kHz 1 kHz to 8 kHz 500 Hz to 8 kHz
@@ V Attack Model Demonstration in Cleanroom
Raw pressure data 125PA 2.95 1.27 9.87 1.14 0.79 78(W) 77(A) 3.15 0.95 8.54 0.98 0.75 85(W) 86(A) 3.45 0.84 6.24 0.87 0.71 97(W) 98(A)
Setra 2.94 1.29 9.86 1.13 0.79 76(W) 78(A) 3.14 0.96 8.57 0.97 0.75 86(W) 87(A) 3.47 0.86 6.22 0.88 0.71 98(W) 99(A)
500PA 2.94 1.29 9.88 1.18 0.79 78(W) 78(A) 3.16 0.98 8.57 0.97 0.75 85(W) 86(A) 3.45 0.84 6.24 0.87 0.71 98(W) 97(A)
Reconstructed 125PA 1.01 2.17 11.38 1.95 0.82 20(W) 21(A) 1.16 1.95 10.04 1.61 0.79 26(W) 27(A) 1.24 1.78 8.78 1.51 0.75 38(W) 39(A)
Setra 1.00 2.18 11.37 1.96 0.82 21(W) 20(A) 1.15 1.96 10.02 1.63 0.79 27(W) 27(A) 1.26 1.80 8.79 1.53 0.75 38(W) 38(A)
500PA 1.00 2.16 11.37 1.97 0.82 21(W) 21(A) 1.17 1.94 10.14 1.60 0.79 26(W) 27(A) 1.24 1.80 8.79 1.53 0.75 36(W) 38(A)
VI Attack Model Evaluation
We demonstrate WaLi at an FDA-compliant anonymous cleanroom in Section V. As it was not permitted to experiment with the HVAC located in the cleanroom for collecting large corpus of pressure data, we prepare a testbed to evaluate WaLi using the same DPS (part# SDP810-125PA), vinyl tubes, and pressure pickup device (see Section V), shown in Fig. 7 (Right). We also verify our work using two more industry-used DPSs – SETRA264 [35] from Setra Systems (Fig. 1 (Right)) and SDP810-500PA [32] (see Appendix C for sensor details).
VI-A Pressure Data and Audio Corpus
@@
We collect pressure data in two different ways. First, we use 20 volunteers (9 male and 11 female) to utter from Wikipedia and collect a total of 8.33 hours of pressure data from each of the three DPSs ($\sim$25 minutes from each volunteer with permission and no ethical concern). Second, as 8.33 hours of data is not enough for proper model training, we use VCTK v0.92 [36], a multi-speaker English corpus, and play the audio at 60 dB through a loudspeaker and collect the corresponding pressure data for evaluation. The dataset contains $\sim$39 hours of audio traces sampled at 48 kHz from 110 individuals (i.e., balanced for males and females, open source and no ethical concern). We downsample the dataset to 8 kHz for evaluation. Each audio clip has a duration ranging from 2s to 7s. We standardize all audio clips to 4s by either zero-padding or silence trimming. The speaker is placed at a 0.5 m distance from one of the pressure ports to play the audio clips. To prevent sound signals from influencing the pressure in the other port, we connect a sampling tube to it, isolated by positioning its opening a meter away. Signals from the sensor are collected as .edf and next converted to .wav format using a custom Python script. The sampling frequency of the DPS is varied in between 500 Hz and 2 kHz. We use sinc interpolation to upsample before the audio reconstruction to ensure that the system input and output have the same shape. Note that in a real case, the speech contents may be different from the spoken ones during the attack phase. Thus, for testing, we use 11 different speakers not present in the training to prove that WaLi reconstructs speech from the victim’s pressure sensor data without using the ground truth audio of the victim. The models are trained offline with an NVIDIA 4090 GPU.
VI-B Comprehensive Evaluation Metrics
To comprehensively evaluate the reconstructed speech, we use six evaluation metrics: log spectral distance (LSD) [26] for spectral distortion, short-time objective intelligibility (STOI) [37] for intelligibility, perceptual evaluation of speech quality (PESQ) [38] for perceived quality, scale-invariant signal-to-distortion ratio (SI-SDR) [39] for overall signal distortion, non-intrusive speech quality assessment - mean opinion score (NISQA-MOS) and word error rate (WER) for perceived quality. We refer to Appendix D for more details on metrics.
VI-C Overall Performance
@@
To intuitively observe the performance of WaLi, we compare among ground truth speech, raw data from the pressure sensor, and reconstructed speech by WaLi (see Fig. 8). The frequency components above 250 Hz are absent in the pressure sensor data (see Fig. 8(b)) as we use a sampling frequency of 0.5 kHz in DPS at this time. The ground truth audio is sampled at 8 kHz (4 kHz bandwidth), shown in Fig. 8(a). From Fig. 8(c), we can see that the high-frequency components are reconstructed up to 4 kHz by WaLi. The reconstructed spectrogram shows a high similarity to the ground truth one.
@@
Moreover, to quantitatively measure WaLi’s performance, we vary the sampling frequency of the pressure sensor on three scales – 500 Hz, 1 kHz, and 2 kHz. We reconstruct the audio to 8 kHz, to evaluate WaLi for high upsampling ratios [26]. Table II shows the results for three industry-used DPSs – SDP810-125PA, SDP810-500PA, and SETRA264 – for combined 8.33-hour HVAC data and VCTK dataset. Table II indicates that WaLi’s performance does not vary for different sensors. WaLi achieves a 2.78x improvement in LSD (i.e., 3.45 vs 1.24), 2.11x increase in NISQA-MOS (i.e., 0.84 vs 1.78), 1.4x increase in SI-SDR (i.e., 6.24 vs 8.78), 1.73x increase in PESQ (i.e., 0.87 vs 1.51), and 1.05x increase in STOI (i.e., 0.71 vs 0.75) between the raw pressure data and the reconstructed audio for 500 Hz sampling for SDP810-125PA.
@@
Please also note that the metrics are more improved for 2-8 kHz compared to 1-8 kHz and 0.5-8 kHz because for 2-8 kHz, the upsampling ratio is around 4, for 1-8 kHz the upsampling ratio is 8, and for 0.5-8 kHz the upsampling ratio is 16. The lower the upsampling ratio, the greater the performance gain for the reconstruction model.
@@
Testing on real-world HVAC dataset: Table II shows results for the combined {8.33-hours real-world HVAC and the VCTK} test dataset. We also evaluate WaLi separately on 20% of 8.33-hour HVAC data and 20% of the VCTK dataset, shown in Table III. We see that WaLi gives similar performance for both the real-world HVAC and the VCTK dataset, indicating WaLi’s effectiveness on real-world HVAC systems.
TABLE III: Evaluating separately on R = Real-world HVAC data, and V = VCTK dataset for 500 Hz to 8 kHz upsampling.
[…]
R V R V R V R V R V
WaLi 1.24 1.23 1.78 1.78 8.78 8.77 1.50 1.51 0.7 0.75
Individual test speaker: In addition to average values, Fig. 9 (Left) shows the individual values of each test speaker for LSD and NISQA-MOS. The individual values indicate that WaLi performs consistently for all unseen test speakers.
Subjective analysis: To assess the perceptual quality of WaLi relative to the raw pressure data, we conduct a listening test with 18 participants using a 5-point Mean Opinion Score (MOS) scale (1 = poor, 5 = excellent). As shown in Fig. 9 (Right), the average MOS of $\sim$2.7 for 500 Hz to 8 kHz and 3.8 for 2 kHz to 8 kHz upsampling indicates that the reconstructed speech achieves a fair to moderate quality relative to the completely unrecognizable raw pressure data (i.e., MOS = $\sim$1). This finding is also supported by the WER discussed next. We refer to Appendix H for the experimental setup for MOS test.
Speech-to-Text accuracy: We evaluate the reconstructed speech using two pretrained models—OpenAI’s Whisper-base [40] and the AssemblyAI’s Speech-to-Text Recognition API [41]—to compute the WER. The resulting WERs (see Table II) are approximately $\sim$20%, $\sim$26%, and $\sim$38% for sampling frequencies of 2 kHz, 1 kHz, and 500 Hz, respectively, substantially lower than the $\sim$78-98% WER observed for the raw pressure data. These findings reinforce that WaLi consistently reconstructs higher perceptual quality speech favored by human listeners, transforming pressure signals that were otherwise completely incomprehensible in their raw form.
Figure 9: (Left) LSD and NISQA-MOS for individual test speakers. (Middle) Comparison between WaLi and VibSpeech [10]. (Right) MOS score.
Comparison with VibSpeech [10]: We also compare between WaLi and VibSpeech [10] in terms of SNR and Mel-Ceptral Distortion (MCD) [10] in Fig. 9 (Middle), which indicates that WaLi performs better (MCD = 3.61 and SNR = 6.3 dB) than VibSpeech (MCD = 3.7 and SNR = 6.1 dB).
Please note that WaLi is tested with different speakers, which are not seen by the model during the training time. Moreover, the VCTK has everyday-use audio clips, which are not limited to a certain specialized vocabulary, such as [24]. Therefore, WaLi can recover moderately intelligible audio with unrestricted vocabulary, irrespective of applications.
TABLE IV: Evaluation of phase reconstruction at non-noisy conditions.
@@
VI-D Phase Reconstruction in Non-Noisy Condition
Previous works used the Griffin-Lim algorithm [13] as a vocoder to reconstruct the phase. To compare the performance of our phase reconstruction in the non-noisy condition, we reconstruct audio from only the magnitude of our model with the Griffin-Lim algorithm, and then compare it with the reconstructed audio from our model. A comparison summary is shown in Table IV for a 500 Hz to 8 kHz reconstruction. Table IV indicates that our proposed complex-valued WaLi is similar/slightly better in phase reconstruction compared to the Griffin-Lim algorithm in the non-noisy condition. However, the performance of WaLi is much better under noisy conditions that is discussed in the next section.
TABLE V: Evaluation of phase reconstruction at noisy conditions separately for R = Real-world HVAC data & D = DCASE data.
@@ VI-D Phase Reconstruction in Non-Noisy Condition
R D R D R D R D R D
Griffin-Lim 1.38 1.40 1.43 1.41 6.17 6.19 1.21 1.23 0.70 0.69
WaLi 1.26 1.28 1.71 1.73 8.14 8.18 1.47 1.49 0.74 0.73
VI-E Phase Reconstruction in Noisy Conditions
DPSs, operating within a low-pressure range of 0-200 Pa and at high sampling frequencies of 0.5-2 kHz, are sensitive to transient noise sources (see Section II-D). Therefore, clean phase reconstruction becomes particularly significant in noisy conditions. To test the performance of phase reconstruction under noisy conditions, we use 80% of the 8.33-hour dataset collected previously from the real-world HVACs to train the model. This dataset has the actual noisy condition of a real-world HVAC system. Moreover, as this real-world HVAC noisy data is not enough to train our model, we also use the DCASE challenge dataset [42] to add transient mechanical noise, such as fans, valves, pump noise, air leaks, tool drops, etc., to the clean VCTK dataset. Noise is randomly added within a range of -15 to 15 dB. Then, we retrain WaLi with a {noisy, clean} pair combined from both the real-world and DCASE datasets. We test WaLi separately using the real-world noisy data collected from the HVAC and DCASE dataset. A comparison summary is shown in Table V for a 500 Hz to 8 kHz reconstruction in noisy conditions. Table V indicates that our proposed complex-valued WaLi is much better in phase reconstruction compared to the Griffin-Lim algorithm in transient noisy conditions for both real-world HVAC and DCASE datasets. The reason behind this is that WaLi handles both magnitude and phase jointly and recovers both clean magnitude and phase under noisy conditions from the clean reference signal using a complex multi-resolution loss function. On the other hand, Griffin-Lim-based models [13] do not learn clean phase reconstruction as these models are magnitude-only models and cannot handle phase jointly.